OpenStudy (anonymous):

What's the square root of 45x^2

OpenStudy (anonymous):

3x square root 5

OpenStudy (anonymous):

\[3x \sqrt{5}\]

OpenStudy (anonymous):

Well since you have a sqr on the x variable it will become just plain old x, then you find the square rt of 45.

OpenStudy (jhannybean):

\[\large \sqrt{45x^2}\]\[\large \sqrt{\color{green}{3 \cdot 3} \cdot 5 \cdot \color{red}{x \cdot x}}\] for every pair of highlighted numbers, 1 of them us removed from the square root. \[\large \color{green}3 \color{red}x \sqrt{5}\]

OpenStudy (tkhunny):

Not quite. \(\sqrt{45x^{2}} = 3|x|\sqrt{5}\). The absolute value is NOT optional unless you know more about x than you are saying.

OpenStudy (anonymous):

\[\root 45x^2= \root 9x^2 * \root 5\]